OpenStudy (anonymous):

Which line is perpendicular to ax+5y=2? a. -5x-ay=2 b. ax-5y=2 c. -ax+5y=2 d. -5x+ay=2

OpenStudy (anonymous):

slope of \(ax+5y=2\) is \(-\frac{a}{5}\) find the choice where the slope is \(\frac{5}{a}\)

OpenStudy (anonymous):

The answer would be A then. Thank you @satellite73!!!

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

\[-5x+ay=2\] has slope \(\frac{5}{a}\)